OpenStudy (anonymous):

What is the slope of the line that passes through the pair of points? (1, 7), (10, 1) A. 3/2 B. -2/3 C. -3/2 D. 2/3

OpenStudy (anonymous):

i think it D

OpenStudy (anonymous):

you could use the equation (y2 - y1)/(x2 - x1) to solve for the slope

OpenStudy (anonymous):

y1-y2/x2-x1

OpenStudy (anonymous):

i mean y2-y1

OpenStudy (anonymous):

the answer would be B

OpenStudy (anonymous):

sorry D* your right

OpenStudy (anonymous):

its B (1 - 7)(10-1) = -6/9=-2/3

OpenStudy (anonymous):

@bballin629 what do you think

OpenStudy (anonymous):

yeah..wow don't know what I was thinking

OpenStudy (anonymous):

guess I did it in my head a little too fast